# Flash Diffusion: Accelerating Any Conditional Diffusion Model for Few Steps Image Generation

### v2

@@ 5.2 Ablation Study
We also assess the influence of the guidance scale used during training to denoise the input samples with the teacher. For this ablation, unlike in Sec. 5.1, we generate samples from the teacher model using a fixed guidance scale $\omega$ set to either $1,3,5,7,10,13$ or $15$. We report the evolution of the FID and CLIP score according to the guidance scale in the bottom left graph in Figure 6. As pictured in the figure and in line with the behavior observed with the teacher, the choice of the guidance scale has a strong impact on the final performance. While the CLIP score measuring prompt adherence tends to increase with the guidance scale, there exists a trade-off with the FID score that tends to eventually increase with the guidance scale resulting in a potential loss of image quality. We represent by the red dot the setting that we propose which consists in uniformly sampling a guidance scale within a given range $\omega\sim\mathcal{U}([\omega_{\min},\omega_{\max}])$ where $\omega_{\min}$ is set to 3 and $\omega_{\max}$ to 13. This approach reveals to be an interesting choice since it seems to allow the student model to be able to generate samples with a wider diversity.
5.3 On the Method’s Versatility
@@
To highlight the versatility of the proposed method, we apply the same approach to diffusion models trained with different conditionings, different backbones, or using adapters [46]. In the following section, all the models are trained on the LAION dataset [63] with samples having an aesthetic score above 6 and re-captioned using CogVLM [71]. The student models are LoRAs [20] and share the same architecture as the teacher.
5.3.1 Backbones’ Study
Table 3: FID and CLIP score computed on the 10k first prompts of MS COCO2014 validation set for SDXL as teacher.
Method NFE FID $\downarrow$ CLIP $\uparrow$
Teacher (SDXL) 40 18.42 0.339
LCM [42] 8 21.73 0.327
Turbo [61] 4 23.69 0.337
SDXL-lightning [33] 4 24.61 0.329
SDXL-lightning-LoRA [33] 4 25.13 0.328
Hyper-SD-LoRA [56] 4 27.76 0.333
Ours 4 21.62 0.327
Table 4: FID and CLIP score computed on the 10k first prompts of MS COCO2014 validation set for Pixart-$\alpha$ as teacher.
Method NFE FID $\downarrow$ CLIP $\uparrow$
Teacher (Pixart-$\alpha$) 40 28.09 0.316
Ours 4 29.30 0.303
@@ 5.3.1 Backbones’ Study
In this section, we illustrate the ability of the proposed method to adapt to another more recent teacher model. To do so, we elect the publicly available SDXL model [50]. We train a LoRA student model (108M trainable parameters) sharing the same architecture as the teacher for 20k iterations on 4 H100-80Gb GPUs (amounting to a total of 176 H100 hours of training) with a batch size of 2 and a learning rate of $1e^{-5}$ together with Adam optimizer [26] for both the student and the discriminator. We provide in Table 4 the FID and CLIP score computed on the 10k first prompts of COCO2014 validation set. We compare the proposed approach to several distillation methods proposed in the literature. The results for the competitors are computed using publicly available checkpoints. Our method can outperform peers in terms FID of 21.62 while maintaining quite good prompt alignment capabilities. In addition, we also provide a visual overview of the generated samples in Figure 7 for the teacher, the trained student model and LoRA-compatible approaches proposed in the literature (LCM [42], SDXL-lightning [33] and Hyper-SD [56]). Teacher samples are generated with a guidance scale of 5. For a fair comparison with competitors, we include prompts used in [33] for this qualitative evaluation. The proposed approach appears to be able to generate samples that are visually closer to the learned teacher distribution. In particular, HyperSD and lightning seem to struggle to generate samples that are realistic despite creating sharp samples. See the appendices for the comprehensive experimental setup used for this section and additional comparison.
(a) Teacher
@@
A photograph of a school bus in a magic forest
A monkey making latte art
A majestic lion stands proudly on a rock, overlooking the vast African savannah
Figure 7: Application of Flash Diffusion to a SDXL. Teacher samples are generated with a guidance scale of 5. The proposed approach is compared to LoRA based competitors and appears to be able to generate samples that are visually closer to the learned teacher distribution. Best viewed zoomed in.
(a) Teacher
(8 NFEs)
(b) Teacher
@@
A whale with a big mouth and a rainbow on its back jumping out of the water
A small cactus with a happy face in the Sahara desert
A close-up of a person with a shaved head, gazing downwards, with a hand resting on their forehead
Figure 8: Application of Flash Diffusion to a DiT-based Diffusion model (Pixart-$\alpha$). The proposed method 4 NFEs generations are compared to the teacher generations using 8 NFEs and 40 NFEs as well as Pixart-LCM [43] with 4 steps. Teacher samples are generated with a guidance scale of 3.
@@
In addition to UNet-based architecture, we propose in this section to apply the proposed method to a DiT denoiser backbone [49]. This appeared important to us since there exist only a few distillation methods that showed to be able to efficiently distill DiT models [43, 77] while these models have demonstrated very promising performances for text-to-image generations [5, 6, 11] or video generation [4]. To do so, we elect the Pixart-$\alpha$ model [5] as teacher. We train a LoRA student model (66.5M trainable parameters) sharing the same architecture as the teacher for 40k iterations on 4 H100-80Gb GPUs (amounting to a total of 188 H100 hours of training) with a batch size of 2 and a learning rate of $1e^{-5}$ together with Adam optimizer [26] for both the student and the discriminator. We compare the student generations using 4 NFEs to the teacher generations using 8 NFEs (4 steps) and 40 NFEs (20 steps) as well as Pixart-LCM [43] using 4 steps in Figure 8. We use a guidance scale of 3 to generate with the teacher. As illustrated in the figure, the proposed method can generate high-quality samples that sometimes seem even more visually appealing than the teacher. Moreover, driven by the adversarial approach the student model trained with our method generates images with more vivid colors and sharper details than LCM. It is noteworthy, that the student model does not loose the capability of the teacher to generate samples that are coherent with the prompt. In addition to those qualitative results, we provide in Table 4 the FID and CLIP score computed on the 10k first prompts of COCO2014 validation set for our model and the teacher. See the appendices for the comprehensive experimental setup and additional samples as well as discussion on the variability of the output samples with respect to the prompt
5.3.2 Conditionings’ Study
@@
(b) Masked
Image
(c) Teacher         (8 NFEs)
(d) Teacher      (40 NFEs)
(e) Ours
(4 NFEs)
(f) LR image
(g) Teacher         (8 NFEs)
(h) Teacher         (40 NFEs)
(i) Ours
(4 NFEs)
Figure 9: Application of Flash Diffusion to an in-house diffusion-based inpainting model (left) and in-house diffusion-based super-resolution model (right). Note that a single step for the teacher involves 2 NFEs since it uses CFG. Best viewed zoomed in.
Inpainting
First, we consider an in-house inpainting model whose backbone is common to SDXL [50]. The model is conditioned on both a masked image, a mask, and a prompt and is trained to reconstruct an input image from a masked one. Again, we initialize a student model with the weights of the teacher and train the student for 20k iterations. The main hyper-parameters are the same as in Sec. 5.1 unless the guidance that was sampled in $\mathcal{U}([5.0,10.0])$ and $K$ is set to 16 to speed up the training. We show some samples in Figure 9 (left) and compare the samples generated by the student model using 4 NFEs to the teacher generations using 4 steps (i.e. 8 NFEs) and 20 steps (i.e. 40 NFEs). In all cases, the models take as input the masked image and an associated prompt to generate the output image. As highlighted in the figure, the proposed method is able to generate samples that are visually close to the teacher generations while using far fewer NFEs demonstrating the ability of the method to adapt to different conditionings and tasks. See the appendices for the comprehensive experimental setup and additional samples.
Super-Resolution
To stress the method in cases where there is no text conditioning, we propose to consider an in-house diffusion model trained to upscale input images by a factor of 4. The model is conditioned on a low-resolution input image and is trained to generate an upscaled image. The degradation process used to artificially create the low-resolution image is similar to the one proposed in [72]. We show some generated samples using either the teacher model with 8 NFEs or 40 NFEs and the student model with 4 NFEs in Figure 9 (right). As highlighted in the figure, the proposed method can generate samples that are visually on par with the teacher generations while using far fewer NFEs. For this use case, tiling can be applied at inference time to generate images of resolution higher than the one used during training ($1024\times 1024$). Nonetheless, the associated computational cost then scales quadratically with the image resolution becoming prohibitively expensive for the teacher. This emphasizes the usefulness of the proposed approach that reduces drastically the number of NFEs required to generate samples and so unlocks the possibility of generating high-resolution images. See the supplementary material for the comprehensive experimental setup and additional samples.
Face-Swapping
(a) Source image
(b) Target image
(c) Teacher
(8 NFEs)
(d) Teacher
(40 NFEs)
(e) Ours
(4 NFEs)
Figure 10: Application of Flash Diffusion to an in-house diffusion-based face-swapping model. Best viewed zoomed in.
In addition to text-to-image, inpainting, and super-resolution, we also propose to apply the proposed method to an in-house face-swapping model. The model is conditioned on a source image and is trained to replace the face of the person in the target image with the one in the source image. In Figure 10, samples obtained using the teacher model with 8 NFEs or 40 NFEs are compared to the student model samples obtained with 4 NFEs. Interestingly, for this task, the proposed method can generate samples that seem to qualitatively outperform the teacher generation quality. The visual inspection of the results seems to indicate that the student appears to better respect the identity of the person in the source image and create more realistic output samples. This again highlights the ability of the method to adapt to different use cases. See the appendices for the comprehensive experimental setup.
Adapters
Finally, we show the compatibility of the proposed approach with adapters. For this experiment, we consider the canny and depth SDXL T2I adapters [46]. In these cases, the student model is trained to output samples conditioned on both a prompt and an additional conditioning given either by the edges of an input image or a depth map. The student model is trained using the proposed method and the same hyper-parameters as in Sec. 5.1 unless the guidance that was sampled in $\mathcal{U}([3.0,7.0])$ and $K$ is set to 16 to speed up the training. For both adapters, we use a conditioning scale of 0.8 to generate the samples with the student model. We show some samples in Figure 11.
(a) Canny T2I Adapters
(b) Depth Map T2I Adapters
Figure 11: Application of Flash Diffusion to canny and depth T2I adapters [46]. Samples are generated using 4 NFEs Best viewed zoomed in.
6 Conclusion
In this paper, we proposed a new versatile, fast, and efficient distillation method for diffusion models. The proposed method relies on the training of a student model to generate samples that are close to the data distribution learned by a teacher model using a combination of a distillation loss, an adversarial loss, and a distribution matching loss. We also proposed to rely on the LoRA method to reduce the number of training parameters and speed up the training process. We evaluated the proposed method on a text-to-image task and showed that it can achieve SOTA results on COCO2014 and COCO2017 datasets. We also stressed and illustrated the versatility of the method by applying it to several tasks (inpainting, super-resolution, face-swapping), different denoiser architectures (UNet, DiT), and adapters where the trained student model was able to produce high-quality samples using only a few numbers of NFEs. Future work would consists in trying to reduce even more the number of NFEs or trying to enhance the quality of the samples by applying Direct Preference Optimization [52, 70] directly to the student model.
Appendix A Experimental Details
A.1 Experimental Setup for Text-to-Image
To compute the FID, we rely on the clean-fid library [48] while we use an OpenCLIP-G backbone [21] to compute the CLIP scores. The models are trained on the LAION dataset [63] where we select samples with aesthetic scores above 6 and re-caption the samples using CogVLM [71].
Flash SD1.5
In this section, we provide the detailed experimental setup used to perform the quantitative evaluation of the model described in Sec. 5.1. For this experiment, we use SD1.5 model as teacher and initialize the student with SD1.5’s weights. The student model is trained for 20k iterations on 2 H100-80Gb GPUs (amounting to 26 H100 hours of training) with a batch size of 4 and a learning rate of $10^{-5}$ for both the student and the discriminator. We use the timestep distribution $\pi(t)$ detailed in Sec. 4.2 with $K=32$ and shift modes every 5000 iterations. We also start with both $\lambda_{\mathrm{adv}}=0$ and $\lambda_{\mathrm{DMD}}=0$ and progressively increase each time we change the timestep distribution so they reach final values set to $0.3$ and $0.7$ respectively. The schedule is $[0,0.1,0.2,0.3]$ for $\lambda_{\mathrm{adv}}$ and $[0,0.3,0.5,0.7]$ for $\lambda_{\mathrm{DMD}}$. The guidance scale $\omega$ used to denoise using the teacher model is uniformly sampled from $[3,13]$. The distillation loss is set to the MSE loss and the GAN loss is set to the LSGAN loss.
When ablating the timesteps distribution, we use the following distributions: $\pi^{\mathrm{uniform}}(t)$, $\pi^{\mathrm{gaussian}}(t)$, $\pi^{\mathrm{sharp}}(t)$ and $\pi^{\mathrm{ours}}(t)$ that are represented in Figure 12.
(a) (a) $\pi^{\mathrm{uniform}}$
(b) (b) $\pi^{\mathrm{gaussian}}$
[Warm-up]Refer to caption [Phase 1]Refer to caption [Phase 2]Refer to caption [ Phase 3]Refer to caption
(c) (c) $\pi^{\mathrm{sharp}}$
[Warm-up]Refer to caption [Phase 1]Refer to caption [Phase 2]Refer to caption [ Phase 3]Refer to caption
(d) (d) $\pi^{\mathrm{ours}}$
Figure 12: Illustration of the timestep distributions used in the ablation study in Sec 5.2
Flash SDXL
In this section, we train a student model sharing the same UNet architecture as SDXL. The model is trained for 20k iterations on 4 H100-80Gb GPUs with a batch size of 2 and a learning rate of $10^{-5}$ for both the student and the discriminator. The student weights are initialized with the teacher’s one. The timestep distribution $\pi(t)$ is detailed in Sec. 4.2 and chosen such that $K=32$. We also shift modes every 5000 iterations. As for SD1.5, we set $\lambda_{\mathrm{adv}}=0$ and $\lambda_{\mathrm{DMD}}=0$ and progressively increase each time we change the timestep distribution so they reach final values set to $0.3$ and $0.7$ respectively. The schedule is $[0,0.1,0.2,0.3]$ for $\lambda_{\mathrm{adv}}$ and $[0,0.3,0.5,0.7]$ for $\lambda_{\mathrm{DMD}}$. We use a guidance scale $\omega$ uniformly sampled from $[3,13]$ with a distillation loss chosen as LPIPS and the GAN loss is set to the LSGAN loss.
Flash Pixart (DiT)
We train a student model sharing the same architecture as the teacher for 40k iterations on 4 H100-80Gb GPUs with a batch size of 2 and a learning rate of $1e^{-5}$ together with Adam optimizer [26] for both the student and the discriminator. The weights of the student model are initialized using the teacher’s. We use the timestep distribution $\pi(t)$ detailed in Sec. 4.2 with $K=16$ and shift modes every 10000 iterations. We also start with both $\lambda_{\mathrm{adv}}=0$ and $\lambda_{\mathrm{DMD}}=0$ and progressively increase each time we change the timestep distribution so they reach final values set to $0.3$ and $0.7$ respectively. The schedule is $[0,0.05,0.1,0.2]$ for $\lambda_{\mathrm{adv}}$ and $[0,0.3,0.5,0.7]$ for $\lambda_{\mathrm{DMD}}$. The guidance scale $\omega$ used to denoise using the teacher model is uniformly sampled from $[2,9]$. The distillation loss is LPIPS loss and the GAN loss is set as the LSGAN loss.
A.2 Experimental Setup for Inpainting
For the inpainting experiment, we use an in-house diffusion-based model whose backbone architecture is similar to the one of SDXL [50] and weights are initialized using the teacher. The student model is trained on 512x512 input image resolution for 20k iterations on 2 H100-80Gb GPUs with a batch size of 4 and a learning rate of $10^{-5}$ for both the student and the discriminator. The timestep distribution $\pi(t)$ is detailed in Sec. 4.2 and chosen with $K=16$. Modes are shifted every 5000 iterations. We again start with both $\lambda_{\mathrm{adv}}=0$ and $\lambda_{\mathrm{DMD}}=0$ and progressively increase each time we change the timestep distribution so they reach final values set to $0.3$ and $0.7$ respectively. The schedule is $[0,0.1,0.2,0.3]$ for $\lambda_{\mathrm{adv}}$ and $[0,0.3,0.5,0.7]$ for $\lambda_{\mathrm{DMD}}$. The guidance scale $\omega$ is uniformly sampled from $[3,13]$. The distillation loss is set as the MSE loss and the GAN loss is set as the LSGAN loss.
A.3 Experimental Setup for Super-Resolution
For the super-resolution experiment, we use an in-house diffusion-based model whose backbone architecture is similar to the one of SDXL [50]. The student model is trained with 256x256 low-resolution images used as conditioning and output 1024x1024 images. The student model is initialized using the teacher’s weights and is trained for 20k iterations on 2 H100-80Gb GPUs with a batch size of 4 and a learning rate of $10^{-5}$ for both the student and the discriminator. We set $K=16$ for $\pi(t)$ and shift modes every 5000 iterations. We start with $\lambda_{\mathrm{adv}}=0$ and $\lambda_{\mathrm{DMD}}=0$ and progressively increase each time we change the timestep distribution so they reach final values set to $0.3$ and $0.7$ respectively. The schedule is $[0,0.1,0.2,0.3]$ for $\lambda_{\mathrm{adv}}$ and $[0,0.3,0.5,0.7]$ for $\lambda_{\mathrm{DMD}}$. The guidance scale $\omega$ used to denoise using the teacher model is uniformly sampled from $[1.2,1.8]$. The distillation loss is set as the MSE loss and the GAN loss is chosen as the LSGAN loss.
A.4 Experimental Setup for Face-Swapping
For the face-swapping experiment, we use an in-house diffusion-based model whose backbone architecture is similar to the one of SD2.2 [57]. The student model is trained on 512x512 input images and target images. We use a face detector to extract the face from the source image and use it as conditioning. The student model is then initialized using the teacher’s weights and is trained for 15k iterations on 2 H100-80Gb GPUs with a batch size of 8 and a learning rate of $10^{-5}$ for both the student and the discriminator. We use the timestep distribution $\pi(t)$ detailed in Sec. 4.2 with $K=16$ and shift modes every 5000 iterations. We also start with both $\lambda_{\mathrm{adv}}=0$ and $\lambda_{\mathrm{DMD}}=0$ and progressively increase each time we change the timestep distribution so they reach final values set to $0.3$ and $0.7$ respectively. The schedule is $[0,0.1,0.2,0.3]$ for $\lambda_{\mathrm{adv}}$ and $[0,0.3,0.5,0.7]$ for $\lambda_{\mathrm{DMD}}$. The guidance scale $\omega$ used to denoise using the teacher model is uniformly sampled from $[2.0,7.0]$. The distillation loss is set as the MSE loss and the GAN loss is chosen as the LSGAN loss.
Appendix B Additional Sampling Results
In this section, we provide additional samples for each task considered in the main paper.
B.1 Flash SDXL
In Figure 13, we provide addition samples enriching the qualitative comparision performed in the main manuscript. Again, to be fair to the competitors, we use some prompts from [33] to generate the samples. As mentioned in the paper, the proposed approach appears to be able to generate samples that are visually closer to the learned teacher distribution.
(a) Teacher
(40 NFEs)
(b) LCM
(4 NFEs)
(c) Lightning
(4 NFEs)
(d) HyperSD
(4 NFEs)
(e) Ours
(4 NFEs)
A pickup truck going up a mountain switchback
A giant wave breaking on a majestic lighthouse
An Asian firefighter with a rugged jawline rushes through the billowing smoke of an autumn blaze
Cute cartoon small cat sitting in a movie theater eating popcorn, watching a movie
A very realistic close up of an old elderly man with green eyes looking straight at the camera, vivid colors
A delicate porcelain teacup sits on a saucer, its surface adorned with intricate blue patterns
Figure 13: Application of Flash Diffusion to a SDXL teacher model. The proposed method 4 NFEs generations are compared to the teacher generations using 40 NFEs as well as LoRA approaches proposed in the literature (LCM [42], SDXL-lightning [33] and Hyper-SD [56]). Teacher samples are generated with a guidance scale of 5. Best viewed zoomed in.
B.2 Flash Pixart (DiT)
In this section, we provide additional samples using the trained student model using a DiT architecture. In Figure 14, we provide a more complete qualitative comparison with respect to LCM and the teacher model while in Figures 15 and 16, we show additional samples using the proposed method. In Figures 17 and 18, we also show the generation variation with respect to two different prompts: A yellow orchid trapped inside an empty bottle of wine and An oil painting portrait of an elegant blond woman with a bowtie and hat. The model appears to be able to generate various samples even with a fixed prompt.
(a) Teacher
(8 NFEs)
(b) Teacher
(40 NFEs)
(c) LCM
(4 NFEs)
(d) Ours
(4 NFEs)
A cute cheetah looking amazed and surprised
A giant wave shoring on big red lighthouse
A raccoon reading a book in a lush forest
A classic turquoise car is parked outside a modern building with curved balconies
A beautiful sunflower in rainy day
A woman in a red traditional outfit wields a sword, poised in an intense stance against a dark background
Figure 14: Application of Flash Diffusion to a DiT-based Diffusion model, namely Pixart-$\alpha$. The proposed method 4 NFEs generations are compared to the teacher generations using 8 NFEs and 40 NFEs as well as Pixart-LCM [43] with 4 steps. Teacher samples are generated with a guidance scale of 3.
(a) A famous professor giraffe in a classroom standing in front of the blackboard teaching
(b) A close up of an old elderly man with green eyes looking straight at the camera
(c) A cute fluffy rabbit pilot walking on a military aircraft carrier, 8k, cinematic
(d) Pirate ship sailing on a sea with the milky way galaxy in the sky and purple glow lights
Figure 15: Application of Flash Diffusion to a DiT-based Diffusion model Pixart-$\alpha$.
(a) A photograph of a woman with headphone coding on a computer, photograph, cinematic, high details, 4k
(b) A super realistic kungfu master panda Japanese style
(c) The scene represents a desert composed of red rock resembling planet Mars, there is a cute robot with big eyes feeling alone, It looks straight to the camera looking for friends
(d) A serving of creamy pasta, adorned with herbs and red pepper flakes, is placed on a white surface, with a striped cloth nearby
Figure 16: Application of Flash Diffusion to a DiT-based Diffusion model Pixart-$\alpha$.
Figure 17: Generation variation for Flash Pixart with the prompt A yellow orchid trapped inside an empty bottle of wine.
Figure 18: Generation variation for Flash Pixart with the prompt An oil painting portrait of an elegant blond woman with a bowtie and hat.
B.3 Flash Inpainting
In Figure 19, we provide additional samples using the trained inpainting student model. We compare the samples generated by the student model using 4 NFEs to the teacher generations using 4 steps (i.e. 8 NFEs) and 20 steps (i.e. 40 NFEs).
(a) Original image
@@ B.3 Flash Inpainting
(c) Teacher
(8 NFEs)
(d) Teacher
(40 NFEs)
(e) Ours
(4 NFEs)
Figure 19: Application of Flash Diffusion to an in-house diffusion-based inpainting model. Best viewed zoomed in.
B.4 Flash Upscaler
In Figure 20, we provide additional samples using the trained super-resolution student model. As in the main paper, the student model is trained to output 1024x1024 images using 256x256 low-resolution images as conditioning. It is compared to the teacher generations using 4 steps (i.e. 8 NFEs) and 20 steps (i.e. 40 NFEs).
(a) LR image
(b) Teacher
(8 NFEs)
(c) Teacher
(40 NFEs)
(d) Ours
(4 NFEs)
Figure 20: Application of Flash Diffusion to an in-house diffusion-based super-resolution model. Best viewed zoomed in.

### v3

@@ Introduction
For this ablation, unlike in the previous sections, we generate samples from the teacher model using a fixed guidance scale $\omega$ set to either $1,3,5,7,10,13$ or $15$. We report the evolution of the FID and CLIP score accordingly in graph (c) in Fig 5. In line with the behavior observed with the teacher, the choice of the guidance scale has a strong impact on the final performance. While the CLIP score measuring prompt adherence tends to increase with the guidance scale, there exists a trade-off with the FID score that eventually increases with the guidance scale resulting in a potential loss of image quality. We represent by the red dot the setting that we propose which consists in uniformly sampling a guidance scale within a given range.
On the Method’s Versatility
@@
To highlight the versatility of the proposed method, we apply the same approach to diffusion models trained with different conditionings, backbones, or adapters (Mou et al. 2024).
Backbones’ Study
@@
In this section, we illustrate the ability of the method to adapt to a SDXL (Podell et al. 2023) teacher model. We provide in Table 1, the FID and CLIP score computed on the 10k first prompts of COCO2014 validation set. We compare the proposed approach to several distillation methods proposed in the literature using publicly available checkpoints. Our method can outperform peers in terms of FID while maintaining quite good prompt alignment capabilities. In addition, we also provide a visual overview of the generated samples in Fig. 6 for the teacher, the trained student model and LoRA-compatible approaches proposed in the literature (LCM (Luo et al. 2023a), SDXL-lightning (Lin, Wang, and Yang 2024) and Hyper-SD (Ren et al. 2024)). Teacher samples are generated with a guidance scale of 5. For a fair comparison with competitors, we include prompts used in (Lin, Wang, and Yang 2024) for this qualitative evaluation. The proposed approach appears to be able to generate samples that are visually closer to the learned teacher distribution. In particular, HyperSD and lightning seem to struggle to generate samples that are realistic despite creating sharp samples. See the appendices for the comprehensive experimental setup and additional comparisons. Additionally, since our student share the same architecture as the teacher, we notice that our approach can be combined with existing LoRAs in a training-free manner. We show at the bottom right of Fig. 7, 4 steps generations for 6 existing SDXL LoRAs directly plugged to our trained Flash SDXL model. We provide additional samples in the appendices.
Model (# NFE) FID $\downarrow$ CLIP $\uparrow$
SDXL (40) 18.4 33.9
LCM (8) 21.7 32.7
Turbo (4) 23.7 33.7
Lightning (4) 24.6 32.9
Lightning (4) 25.1 32.8
HyperSD (4) 27.8 33.3
Ours (4) 21.6 32.7
LoRAs
Model (# NFE) FID $\downarrow$ CLIP $\uparrow$
Pixart (40) 28.1 31.6
Ours (4) 29.3 30.3
Model (# NFE) FID $\downarrow$ CLIP $\uparrow$
SD3 (40) 24.4 33.5
Ours (4) 27.5 32.8
Table 1: FID and CLIP score on 10k samples of COCO2014 validation set for SDXL, Pixart-$\alpha$ and SD3 teacher.
(a) SDXL
@@
(f) Pixart-$\alpha$
@@
(i) SD 3
(40 NFEs)
(j) Ours
(4 NFEs)
Figure 6: From left to right: Application of Flash Diffusion to SDXL (UNet), Pixart-$\alpha$ (DiT) and Stable Diffusion 3 (MMDiT) teachers. Teacher samples are generated with a guidance scale of 5, 3, and 5 respectively. The proposed approach is compared to LoRA based competitors and appears to be able to generate samples that are visually closer to the learned teacher distribution. Best viewed zoomed in. Additional samples are provided in the appendices.
@@
In this section, we propose to apply the proposed method to a DiT denoiser backbone (Peebles and Xie 2023) using Pixart-$\alpha$ (Chen et al. 2023) as teacher. We compare the student generations using 4 NFEs to the teacher generations using 40 NFEs (20 steps) as well as Pixart-LCM (Luo et al. 2023b) in Fig. 6 and provide metrics in Table 1. The proposed method can generate high-quality samples that sometimes seem even more visually appealing than the teacher. Moreover, driven by the adversarial approach the student model trained with our method generates images with more vivid colors and sharper details than LCM. It is noteworthy that the student model does not lose the capability of the teacher to generate samples that are coherent with the prompt. In addition, we provide in Table 1 FID and CLIP scores computed on the 10k first prompts of COCO2014 validation set for our model and the teacher. See the appendices for the comprehensive experimental setup and additional samples as well as discussion on the variability of the output samples with respect to the prompt.
Flash SD3 (MMDiT)
Finally, we also show the compatibility of our approach with the recently propose MMDiT architecture of Stable Diffusion 3 (Esser et al. 2024). The method is again able to successfully distil the teacher model and generate samples in only 4 NFEs. We train a 90.4M parameter LoRA model with a batch size of 2 and a learning rate of $1e^{-5}$ together with Adam optimizer (Kingma and Ba 2014) for both the student and the discriminator. We provide in Fig. 6 samples generated with the teacher model and our method and quantitative results in Table. 1.
Conditionings’ Study
Inpainting, Super-Resolution and Face-Swapping
In this section, we consider 1) an in-house inpainting diffusion model conditioned on both a masked image, a mask, and a prompt, 2) a super-resolution model trained to upscale input images by a factor of 4 and 3) a face-swapping model conditioned on a source image and trained to replace the face of the person in the target image with the one in the source image. We show some samples in Fig. 7 using either our student model using 4 NFEs or the teacher generations using 4 steps (i.e. 8 NFEs) and 20 steps. As highlighted in the figure, the proposed method is able to generate samples that are visually close to the teacher generations while using far fewer NFEs demonstrating the ability of the method to adapt to different conditionings and tasks. See the appendices for the comprehensive experimental setup and additional samples.
Adapters
We show the compatibility of the proposed approach with T2I adapters (Mou et al. 2024). In this case, the student model is trained to output samples conditioned on both a prompt and an additional conditioning given either with edges or a depth map. Samples are shown in Fig. 7.
Conclusion
In this paper, we proposed a new versatile, fast, and efficient distillation method for diffusion models. The proposed method relies on the training of a student model to generate samples that are close to the data distribution learned by a teacher model using a combination of a distillation loss, an adversarial loss, and a distribution matching loss. We also proposed to rely on the LoRA method to reduce the number of training parameters and speed up the training process. We evaluated the proposed method on a text-to-image task and showed that it can achieve SOTA results on COCO2014 and COCO2017 datasets. We also stressed and illustrated the versatility of the method by applying it to several tasks (inpainting, super-resolution, face-swapping), different denoiser architectures (UNet, DiT, MMDiT), and adapters where the trained student model was able to produce high-quality samples using only a few number of NFEs. Future work would consists in trying to reduce even more the number of NFEs or trying to enhance the quality of the samples by applying Direct Preference Optimization (Rafailov et al. 2024; Wallace et al. 2023) directly to the student model.
@@ Conclusion
(c) Ref. (8 NFE)
(d) Ref. (40 NFE)
(e) Ours (4 NFE)
(p) Source image
(q) Target image
(r) Ref. (8 NFE)
(s) Ref. (40 NFE)
(t) Ours (4 NFEs)
(ae) LR image
(af) Ref. (8 NFEs)
(ag) Ref. (40 NFEs)
(ah) Ours (4 NFEs)
(am) T2I Adapters
(an) Training-free LoRAs compatibility
Figure 7: From top to bottom: Flash Diffusion applied to 1) an inpainting model, 2) a face-swapping model and 3) a super-resolution model as well as T2I adapters. At the bottom right, we show the 4 steps generations from 6 different LoRAs directly applied on top of Flash SDXL (no training needed).